Find the eigenvalues and eigenvectors of the matrices A =

"

1 2

−8 4#

and B =

"

a b

−b a#

.


1
Expert's answer
2021-06-28T16:32:14-0400

Solution.

1.


"A=\begin{pmatrix}
 1 & 2\\
 -8& 4
\end{pmatrix}"

Solve equation for finding eigenvalues:

"|A-\lambda I|=0"


"\begin{vmatrix}
 1- \lambda& 2 \\
 -8 & 4-\lambda
\end{vmatrix}=\lambda^2-5\lambda+20."

From here we have eigenvalues"\lambda_1=\frac{5-\sqrt{55}i}{2}, \lambda_2=\frac{5+\sqrt{55}i}{2}."

Solve equation for finding the first eigenvector

"(A-\lambda_1 I)v_1=0."

Hence


"\begin{pmatrix}
 \frac{-3+\sqrt{55}i}{2} & 2 \\
 -8 & \frac{3+\sqrt{55}i}{2}
\end{pmatrix}\to
\begin{pmatrix}
 1 & \frac{-3-\sqrt{55}i}{16}\\
 0& 0
\end{pmatrix}"

then "x_2=x_2, x_1=\frac{3+\sqrt{55}i}{16}."

So, "v_1=\begin{pmatrix}
 \frac{3+\sqrt{55}i}{16} \\
 1
\end{pmatrix}."

Solve equation for finding the second eigenvector

"(A-\lambda_2 I)v_2=0."

Hence


"\begin{pmatrix}
 \frac{-3-\sqrt{55}i}{2} & 2 \\
 -8 & \frac{3-\sqrt{55}i}{2}
\end{pmatrix}\to
\begin{pmatrix}
 1 & \frac{-3+\sqrt{55}i}{16}\\
 0& 0
\end{pmatrix}"

then "x_2=x_2, x_1=\frac{3-\sqrt{55}i}{16}."

So, "v_2=\begin{pmatrix}
 \frac{3-\sqrt{55}i}{16} \\
 1
\end{pmatrix}."

2.


"B=\begin{pmatrix}
 a& b\\
 -b& a
\end{pmatrix}"

Solve equation for finding eigenvalues:

"|B-\lambda I|=0"


"\begin{vmatrix}
 a- \lambda& b \\
 -b & a-\lambda
\end{vmatrix}=a^2-\lambda^2+b^2."

From here we have eigenvalues"\lambda_1=\sqrt{a^2+b^2}, \lambda_2=-\sqrt{a^2+b^2}."

Solve equation for finding the first eigenvector

"(B-\lambda_1 I)v_1=0."

Hence

"\begin{pmatrix}
 a- \sqrt{a^2+b^2}& b\\
 -b & a-\sqrt{a^2+b^2}
\end{pmatrix}\to
\begin{pmatrix}
 b & -a-\sqrt{a^2+b^2}\\
 0& 2b\sqrt{a^2+b^2}
\end{pmatrix}" then "x_2=2b\sqrt{a^2+b^2},
x_1=2a\sqrt{a^2+b^2}+2(a^2+b^2)."

So, "v_1=\begin{pmatrix}
 2a\sqrt{a^2+b^2}+2(a^2+b^2)\\
 2b\sqrt{a^2+b^2}
\end{pmatrix}."

Solve equation for finding the second eigenvector

"(B-\lambda_2 I)v_2=0."

Hence

"\begin{pmatrix}
 a+\sqrt{a^2+b^2}& b\\
 -b & a+\sqrt{a^2+b^2}
\end{pmatrix}\to
\begin{pmatrix}
 -b & a+\sqrt{a^2+b^2}\\
 0& -2b\sqrt{a^2+b^2}
\end{pmatrix}" then "x_2=-2b\sqrt{a^2+b^2},
x_1=2a\sqrt{a^2+b^2}-2(a^2+b^2)."

So, "v_2=\begin{pmatrix}
 2a\sqrt{a^2+b^2}-2(a^2+b^2)\\
 -2b\sqrt{a^2+b^2}
\end{pmatrix}."